I have a balloon that can hold 100 liters of air.if I blow up this balloon with 3 moles of oxygen gas at a pressure of 1 atmosphere, what is the temperature of the ballon



1
Expert's answer
2022-03-10T10:45:26-0500

According to the ideal gas law,


"T=\frac{PV}{nR}=\frac{1\ atm\times100\ L}{3\ mol\times0.08206\frac{L\cdot{atm}}{mol\cdot{K}}}=406\ K=406-273=133\degree{C}"


Answer: 133oC